OpenStudy (anonymous):

Find a formula for the inverse of the function: y=ln(x+3)

OpenStudy (across):

Let \(y=\ln(x+3)\). Then, to find its inverse, let \(x=\ln(y+3)\) and solve for \(y\). Can you do that?

OpenStudy (anonymous):

x = lny + ln3 x - ln3 = lny

OpenStudy (pradipgr817):

the terms on the right side of the equation will add like you did only when they multiply i.e. ln(yx3) not for this case.

OpenStudy (anonymous):

I'm not sure how to do it otherwise. I don't think you could do: y/(x+3)=ln

OpenStudy (pradipgr817):

you can write \[e ^{x}=y+3\] I think this might help.

OpenStudy (across):

As @pradipgr817 implied, \(e^{\ln x}=\ln e^x=x\).